please help me with this question on probablity.
A bag Cantains 6 White and 4 black balls, two
balls are drawm at random. Find the probability
that they are one of the same colour:-
Thank you..

Answer by stanbon(48568)

A bag Cantains 6 White and 4 black balls, two
balls are drawm at random. Find the probability
that they are one of the same colour:-
--------------
P(2white or 2black) = P(2white) + P(2black) = 6C2/10C2 + 4C2/10C2
= (6C2 + 4C2)/(10C2) = (15 + 6)/45 = 21/45
=============================================
Cheers,
Stan H.